Clyd's age is 5 years more than twice Bonnie's age. If the product of their ages is 105, how old is Clyde?
Answer by stanbon(75887)

Clyde's age is 5 years more than twice Bonnie's age.
---
C = 2B + 5
B = (C-5)/2
---------------------
If the product of their ages is 105, how old is Clyde?
CB = 105
---
Substitute for "B" and solve for "C":
C[(C-5)/2] = 105
C^2-5C-210 = 0
C = 17.21 years
Note: I graphed it and found this answer.
You could use the quadratic formula.
====================
Cheers,
Stan H.
---